At what net rate (in kW) does heat radiate from a 285 m2 black roof on a night when the roof’s temperature is 25.7ºC and the surrounding temperature is 12.1ºC ? The emissivity of the roof is 0.781.
Answer #1

By Stefan Boltzmanin law, the rate of heat enagy tranffered if a = Aer ( 4 TM T= 12.1+273.15 = 285.25k To = 25.7+273.15:298-8